write an equation of the line that passes through each point with the given slope. (-3, -4), m=3
Answer by nerdybill(7384)

write an equation of the line that passes through each point with the given slope. (-3, -4), m=3
.
Plug given info into the "point-slope" form of a line:
y - y1 = m(x - x1)
y - (-4) = 3(x - (-3))
y + 4 = 3(x + 3)
y + 4 = 3x + 9
y = 3x + 5 (this is what they're looking for in "slope-intercept" form)